# Can We Verify Step by Step for Incorrect Answer Detection?

### v1

@@ 4.1 PDS substantially outperforms ADS
As shown in Table 1, across all 45 subsets in our R2PE, PDS consistently improves ADS performance, yielding an average improvement of $5.10\%$ in terms of F1. Although ADS demonstrates greater precision, it suffers from significantly lower recall, indicating its propensity to overlook numerous examples with $L=F$. In contrast, PDS adeptly balances precision and recall, culminating in superior overall performance relative to ADS (Table 14 and 15). This trend can be observed from Figure 3.
PDS demonstrates a stronger capability to verify the accuracy of the predictions for free-form questions than ADS, with notable average improvements of $7.00\%$ in HotpotQA and $9.80\%$ in 2WikiMultihop. This superior performance can be attributed to the process-oriented verification approach employed by PDS. By addressing the challenges associated with indirectly assessing answer consistency in free-form questions, PDS proves to be more effective in ensuring accuracy. It is important to note that the average improvement metrics provided in Table 1 serve primarily as an indicator of the extent to which PDS enhances performance over ADS. However, a direct comparison of F1 scores between different subsets within R2PE is not recommended. This is because the efficacy of the F1 score as a comparative metric is limited by variations in the ratio of TRUE and FALSE instances across different subsets.
@@ 5 Related Work
5.1 Extensions of Chain-of-Thought
Chain-of-Thought (CoT) (Wei et al., 2022) encourages LLMs to detail their reasoning in natural language steps within few-shot examples. Wang et al. (2022) advance CoT by substituting greedy decoding with sampling techniques. Furthermore, various strategies (Chen et al., 2022; Gao et al., 2023; Zhang et al., 2023a) resort to prompting LLMs to express intermediate steps as Python codes. Given the sensitivity of few-shot learning to examplars, there has been an influx of research aiming to source high-quality demonstrations, from automatic annotations (Kojima et al., 2022; Zhang et al., 2022; Fu et al., 2022; Wan et al., 2023; Xu et al., 2023; Wang et al., 2023; Zhao et al., 2023b; Crispino et al., 2023) to hybrid approaches that blend manual design and automatic construction (Shao et al., 2023; Shum et al., 2023; Diao et al., 2023; Zou et al., 2023). For example, zero-shot-CoT (Kojima et al., 2022) proposes a bifurcated prompting approach; Fu et al. (2022) recommend selecting questions that require more reasoning steps while Diao et al. (2023) suggest selecting the most uncertain questions as demonstrations.
To enhance LLM reasoning across various tasks, novel prompt adaptations (Yao et al., 2022, 2023; Besta et al., 2023; Ding et al., 2023; Zheng et al., 2023; Zhang et al., 2023b) have been developed. ReAct (Yao et al., 2022) synergize reasoning with decision-making actions, while structural variants (Yao et al., 2023; Besta et al., 2023; Ding et al., 2023) tackle tasks requiring elaborate exploration and strategic planning. For complex mathematical problems and knowledge-intensive tasks, PHP (Zheng et al., 2023), cumulative reasoning (Zhang et al., 2023b), and verify-and-edit (VE) (Zhao et al., 2023a) have been crafted. Li et al. (2023) further extend VE framework by dynamically incorporating diverse knowledge sources.
Most prior research focuses predominantly on improving final task performance in terms of metrics, such as accuracy or exact match rates. Although a subset of them offers error analysis or failure cases, none of them extensively explores the relationship between the reasoning process and the resultant predictions. To the best of our knowledge, our work is the first to meticulously investigate whether the veracity of final predictions can be predicted by assessing the reasoning chains.
5.2 Reasoning Chain Quality Evaluation
Analyzing rationales is crucial for understanding the capabilities, limitations, and identifying refinement opportunities (Golovneva et al., 2022; Prasad et al., 2023; He et al., 2023). Methods to evaluate these chains fall into two categories: reference-based, which compares generated rationales to human-annotated "gold standards" (Clinciu et al., 2021; Welleck et al., 2022; Saparov and He, 2022), and reference-free approaches, like ROSCOE (Golovneva et al., 2022), which assess rationales using specific metrics without human-created references. Further advancements include fine-tuning and prompting techniques by Prasad et al. (2023) and He et al. (2023), respectively. Although Prasad et al. (2023); He et al. (2023) show that high-quality reasoning pathways can improve task performance, the direct link between reasoning chain quality and prediction correctness remains unexplored. He et al. (2023) qualitatively analyze reasoning chains based on prediction accuracy, but our work presents a comprehensive quantitative analysis to investigate this connection.
@@ A.2 Queried LLMs
We engaged six different LLMs to collect responses across various tasks to ensure generalizability.
@@
A.3 Extraction & Manual Cleaning
@@
LLMs
GSM8K
StrategyQA
Play
Physics
text-davinci-003
Therefore, the answer is …
GPT-3.5-turbo-instruct
Therefore, …; So, the answer is …
Therefore, the answer is …
Therefore, the answer is …; Therefore, the correct answer is …
GPT-3.5-turbo
So, …; Therefore, …
Therefore, …; So, the answer is …
Therefore, the answer is …
Therefore, the answer is …
Gemini Pro
So, …; Therefore, …; Answer: …; The answer: …
Therefore, …
Therefore, …; …, the answer is …
Therefore, …
Claude-2
Therefore, …
Therefore, …; So in summary, …; So in conclusion, …; So the answer is …
Therefore, …; …, the answer is …
Therefore, the answer is …; So the answer is …; So the formula required is …
Mixtral-8x7b
Therefore, …; So, …
So, the answer is …; Therefore, the answer is …
Therefore, …; In conclusion, …; So, the answer is …; Thus, the answer is …
Answer: …; Therefore, …; So the answer is …; So, the answer is …; The final answer is …
mistral-medium
Therefore, …; So, …
Answer: …; Therefore, …
Answer: …; Therefore, …
Therefore, …; So, the answer is …; Answer: …
@@
In a manner akin to the StrategyQA dataset, each response within the Play dataset is anticipated to be classified as either ’different’ or ’same’. To ensure uniformity, we employed an identical pipeline for the extraction of these answers. And 16 questions are filtered in the (Play from Gemini Pro). The identified answer trigger words for the Play dataset and the number of examples that are manually corrected are listed in Table 6 and  7 correspondingly.
@@ Appendix B More Examples of R2PE
Data Field Name
Description
Example
Question $Q$
The question or claim that is used to query LLM from the original dataset.
What was the nationality of the actor playing a Russian hostage taker in "Hostage for a Day"?
Dataset Name
Which dataset $Q$ is from?
HotpotQA
LLM Name
The LLM to generate responses for $Q$.
GPT-4-0314
Response
The reponses generated by "LLM Name" to answer Question $Q$, containing five responses with reasoning chains and answers $R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})$
First, “Hostage for a Day" is a 1994 TV movie directed by John Candy. Second, the Russian hostage-taker character, Boris, is portrayed by George Wendt, an American actor. The answer is American. …
Final Output $a$
The final answer by majority voting.
American
Ground-truth
The ground-truth answer or answer set provided for question. $Q$.
Canadian
Label $L$
label $L\in\{T,F\}$ to indicate whether the generated answer $a$ matches the ground truth.
F
@@ C.2 Precision and Recall
In Table 1, it is evident that PDS consistently surpasses ADS regarding F1 score. To further elucidate these findings, we also present precision and recall as supplementary metrics in Table 14 and 15. While the ADS exhibits higher precision in most cases, the PDS significantly enhances recall across all subgroups. On average, PDS reduces precision by $5.72\%$ and increases recall by $15.68\%$, suggesting a better balance between precision and recall, culminating in an enhanced F1 score. From practical respect, where the detection of all incorrect instances is paramount, even at the cost of occasionally misidentifying correct answers as potentially incorrect, a higher recall is preferable due to its emphasis on comprehensive falsehood detection.
@@ C.3 About the Threshold
As discussed in Section 2, for subsets of the R2PE benchmark with numeric or free-form answer formats, we set the threshold $H=2.5$ for ADS and $H=0$ for PDS. In contrast, for discriminative tasks, the threshold is set to $4.5$ for ADS and $0.4$ for PDS, respectively.
The reason for setting $H=2.5$ for ADS in the context of numeric or free-form responses is that, with $n=5$, an ADS less than 2.5 indicates that the majority cannot agree on the same answers. For PDS, since $PPSS(A,B)<0$ signifies contradictions between A and B and based on Equation (1), $PDS<0$ suggests the presence of conflicting information or unagreable answers.
For discriminative tasks, where the answers are selected from a predetermined set, the threshold of ADS increases to $4.5$. This adjustment accommodates the structured nature of these tasks. Similarly, from Equation (1), we should adjust the threshold for PDS to $0.4$ accordingly.
@@ C.4 HaloCheck and SelfCheckNLI
HaloCheck (Elaraby et al., 2023) and SelfCheckNLI (Manakul et al., 2023) are two hallucination detection methods for open-ended genereation, which can serve as alternative implementations of PSS. HaloCheck uses the same formula of PPSS as PDS-avg, albeit calculating the average over only a subset of all response pairs: $PSS_{HaloCheck}=mean_{i<j}PPSS(R_{i},R_{j})$. It is important to note that $PPSS(A,B)$ is not symmetric with respect to A and B, inrtoducing a nuanced difference in results. In contrast, SelfCheckNLI adopts a distinct PPSS implementation, leveraging DeBERTa-v3-large (He et al., 2021) finetuned to MNLI (Williams et al., 2018). Additionally, SelfCheckNLI also alters the PSS formulation $PSS_{SelfCheckNLI}=mean_{j}PPSS(R,R_{j})$, where $R$ represents a randomly selected response from the set of highly agreeable answers. This method diverges from the full utilization of all responses by focusing the evaluation on the factuality of a single generated response, as highlighted by Elaraby et al. (2023). This deviation from utilizing the collective information present in all responses leads to a notable performance decrease when replacing our PDS implementation with that of PDS-SelfCheckNLI, as evidenced by the results in Table 2.
@@ Appendix D Combining with verify-and-edit
To edit the rationale $r=s_{1},...,s_{n}$, we first generate one question $q_{i}$ for each sentence $s_{i}$ via prompting. The question $q_{i}$ is to verify the sentence $s_{i}$ in the rationale $r$ and is used as a query to retrieve the relevant context $c_{i}$. $c_{i}$ is composed of the top 3 sentences most similar to query $q_{i}$ ranked by the pre-trained Sentence BERT model (Reimers and Gurevych, 2019) in the retrieved documents. Based on the retrieved context $c_{i}$, the answer for verifying question $q_{i}$ is generated, denoted by $s^{\prime}_{i}$. Finally, we have a new rationale $r^{\prime}=s^{\prime}_{1},...,s^{\prime}_{n}$ that is edited leveraging external knowledge. The prompts for verifying question generation and answering again are provided in Table 16,  17, and  18. Specifically, we use the orical setting from Zhao et al. (2023a) to counteract the influence of retriever performance.
@@ Appendix F Prompts
In our study, we employ modified versions of the original CoT prompts (Wei et al., 2022) for both the GSM8K and StrategyQA datasets. As depicted in Table 21 and 22, the experiments were conducted using setups of 6-shot and 4-shot respectively.
For the Play and Physics, the prompts were meticulously crafted by ourselves. The details of these custom-designed prompts are presented in Table 23 and 24.

### v2

@@ 4.1 PDS Substantially Outperforms ADS
As shown in Table 1, across all 45 subsets in our R2PE, PDS consistently improves ADS performance, yielding an average improvement of $5.10\%$ in terms of F1. Although ADS demonstrates greater precision, it suffers from significantly lower recall, indicating its propensity to overlook numerous examples with $L=F$. In contrast, PDS adeptly balances precision and recall, culminating in superior overall performance relative to ADS (Table 15 and 16). From Table 2, PDS improves ADS in terms of AUC-PR in the majority of cases and yields comparable results (less than $1\%$ decrease) in the remaining ones. On average, PDS results in $3.01\%$ absolute improvement in AUC-PR. Notably, ADS is a fairly good baseline on GSM8K, MATH, HotpotQA, and, 2WikiMultihop to verify the correctness of answers given by CoT self-consistency, which is a widely adopted approach to improve reasoning abilities of LLMs. Our findings demonstrate that PDS can further augment the performance of ADS. For discriminative tasks (StrategyQA, Play, Physics, and Play), ADS has relatively low F1 scores and AUC-PR. This may be attributed to the fixed set of answers (e.g., yes or no) in these tasks, which inherently exhibit higher consistency, making verification based solely on answer agreement challenging. In contrast, PDS can relieve this issue by evaluating reasoning processes.
PDS demonstrates a stronger capability to verify the accuracy of the predictions for free-form questions than ADS, with notable average improvements of $7.00\%$ in HotpotQA and $9.80\%$ in 2WikiMultihop. This superior performance can be attributed to the process-oriented verification approach employed by PDS. By addressing the challenges associated with indirectly assessing answer consistency in free-form questions, PDS proves to be more effective in ensuring accuracy. It is important to note that the average improvement metrics provided in Table 1 serve primarily as an indicator of the extent to which PDS enhances performance over ADS. However, a direct comparison of F1 scores and AUC-PR between different subsets within R2PE is not recommended. This is because the efficacy of the F1 score as a comparative metric is limited by variations in the ratio of TRUE and FALSE instances across different subsets.
@@ 5 Related Work
Extensions of CoT. CoT (Wei et al., 2022) encourages LLMs to detail their reasoning steps with few-shot examples. Wang et al. (2022) advance CoT by substituting greedy decoding with sampling techniques. Moreover, various strategies (Chen et al., 2022; Gao et al., 2023; Zhang et al., 2023a) prompt LLMs to express rationales as Python codes. The search for high-quality exemplars has led to both automated (Kojima et al., 2022; Zhang et al., 2022; Fu et al., 2022; Wan et al., 2023; Xu et al., 2023; Wang et al., 2023; Zhao et al., 2023b; Crispino et al., 2023) and hybrid (Shao et al., 2023; Shum et al., 2023; Diao et al., 2023; Zou et al., 2023) demonstration selection. To enhance LLM reasoning across various tasks, novel prompt adaptations (Yao et al., 2022, 2023; Besta et al., 2023; Ding et al., 2023; Zheng et al., 2023; Zhang et al., 2023b) have been developed. For complex mathematical problems and knowledge-intensive tasks, PHP (Zheng et al., 2023), cumulative reasoning (Zhang et al., 2023b), and verify-and-edit (VE) (Zhao et al., 2023a) have been crafted. Li et al. (2023) further extend VE by incorporating diverse knowledge sources. While most research prioritizes task performance metrics, our study poineers the exploration of the connection between reasoning processes and outcomes, aiming to verify final predictions by analyzing reasoning chains.
Rationale Quality Evaluation. Rationale analysis is key for understanding AI performance and limitations(Golovneva et al., 2022; Prasad et al., 2023; He et al., 2023). Rationale evaluation methods are either reference-based, comparing against a gold standard (Clinciu et al., 2021; Welleck et al., 2022; Saparov and He, 2022), or reference-free, using metrics like ROSCOE (Golovneva et al., 2022). While they link high-quality reasoning to better task performance, the direct impact of reasoning quality on prediction accuracy is under-researched. Our study fills this gap with a detailed quantitative analysis of this relationship.
Uncertainty Estimation. A range of methods measure LLM uncertainty by token probabilities (Kuhn et al., 2023; Malinin and Gales, 2020; Kadavath et al., 2022; Fomicheva et al., 2020); however, they are inapplicable to R2PE, which does not furnish probabilities for responses. Moreover, they concentrate on short answer sequences, aligning more closely with the standard prompting (Brown et al., 2020). In contrast, our work verifies the answer veracity of CoT, which encompasses both the reasoning chains and answer sequences. Different from Kuhn et al. (2023), the entailment model in our work is used to measure the informational similarities between two responses, where the problem setting and detailed implementation are distinct.
@@ A.2 Queried LLMs
We engaged six different LLMs to collect responses across various tasks to ensure generalizability. Specifically, we set the temperature to 0.7, $n=5$ for all datasets and ’topK’ to be 40, ’topP’ to be 0.95 in the Gemini Pro API.
@@
A.3 Extraction and Manual Cleaning
@@
LLMs GSM8K StrategyQA Play Physics
text-davinci-003 Therefore, the answer is …
GPT-3.5-turbo-instruct Therefore, …; So, the answer is … Therefore, the answer is … Therefore, the answer is …; Therefore, the correct answer is …
GPT-3.5-turbo So, …; Therefore, … Therefore, …; So, the answer is … Therefore, the answer is … Therefore, the answer is …
Gemini Pro So, …; Therefore, …; Answer: …; The answer: … Therefore, … Therefore, …; …, the answer is … Therefore, …
Claude-2 Therefore, … Therefore, …; So in summary, …; So in conclusion, …; So the answer is … Therefore, …; …, the answer is … Therefore, the answer is …; So the answer is …; So the formula required is …
Mixtral-8x7b Therefore, …; So, … So, the answer is …; Therefore, the answer is … Therefore, …; In conclusion, …; So, the answer is …; Thus, the answer is … Answer: …; Therefore, …; So the answer is …; So, the answer is …; The final answer is …
mistral-medium Therefore, …; So, … Answer: …; Therefore, … Answer: …; Therefore, … Therefore, …; So, the answer is …; Answer: …
@@
In a manner akin to the StrategyQA dataset, each response within the Play dataset is anticipated to be classified as either ’different’ or ’same’. To ensure uniformity, we employed an identical pipeline for the extraction of these answers. And 16 questions are filtered in the (Play from Gemini Pro). The identified answer trigger words for the Play dataset and the number of examples that are manually corrected are listed in Table 7 and  8 correspondingly.
@@ Appendix B More Examples of R2PE
Data Field Name Description Example
Question $Q$ The question or claim that is used to query LLM from the original dataset. What was the nationality of the actor playing a Russian hostage taker in "Hostage for a Day"?
Dataset Name Which dataset $Q$ is from? HotpotQA
LLM Name The LLM to generate responses for $Q$. GPT-4-0314
Response The reponses generated by "LLM Name" to answer Question $Q$, containing five responses with reasoning chains and answers $R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})$ First, “Hostage for a Day" is a 1994 TV movie directed by John Candy. Second, the Russian hostage-taker character, Boris, is portrayed by George Wendt, an American actor. The answer is American. …
Final Output $a$ The final answer by majority voting. American
Ground-truth The ground-truth answer or answer set provided for question. $Q$. Canadian
Label $L$ label $L\in\{T,F\}$ to indicate whether the generated answer $a$ matches the ground truth. F
@@ C.2 Precision and Recall
In Table 1, it is evident that PDS consistently surpasses ADS regarding F1 score. To further elucidate these findings, we also present precision and recall as supplementary metrics in Table 15 and 16. While the ADS exhibits higher precision in most cases, the PDS significantly enhances recall across all subgroups. On average, PDS reduces precision by $5.72\%$ and increases recall by $15.68\%$, suggesting a better balance between precision and recall, culminating in an enhanced F1 score. From practical respect, where the detection of all incorrect instances is paramount, even at the cost of occasionally misidentifying correct answers as potentially incorrect, a higher recall is preferable due to its emphasis on comprehensive falsehood detection.
@@ C.3 About the Threshold
As discussed in Section 3.3, we set the threshold $H=2.5$ for ADS and $H=0$ for PDS and there is no need to tune the threshold $H$. We assume it as an advantage, given that held-out data for threshold adjustment may not exist in practical settings. For example, if a user asks a question and wants to know whether the answer returned from an LLM is correct or not. In this case, there is no held-out data and a task-agnostic threshold should be preferable.
Following Li et al. (2023), we raise the threshold $H$ for discriminative tasks where the answers are from a fixed set and often have greater answer agreement among responses. Such tasks include binary (yes/no) choices or multiple-choice queries. More concretely, the threshold $H$ for ADS is set to be 4.5 and $0.4$ for PDS for subsets in R2PE associated with StrategyQA, Play, Physics, and FEVER. Despite this change, our thresholds are still not task-specific. That is because we can only in advance whether a query is a discriminative one. In our previous example, if the user asks a yes/no or multiple choice question, we set the threshold to $4.5$ for ADS and $0.4$ for PDS, respectively. Otherwise, we set the threshold to $2.5$ for ADS and 00 for PDS.
The reason for setting $H=2.5$ for ADS in the context of numeric or free-form responses is that, with $n=5$, an ADS less than 2.5 indicates that the majority cannot agree on the same answers. For PDS, since $PPSS(A,B)<0$ signifies contradictions between A and B and based on Equation (1), $PDS<0$ suggests the presence of conflicting information or unagreeable answers.
For discriminative tasks, where the answers are selected from a pre-determined set, the threshold of ADS increases to $4.5$. This adjustment accommodates the structured nature of these tasks. Similarly, from Equation (1), we should adjust the threshold for PDS to $0.4$ accordingly.
@@
We also show the effect of varying the threshold of PDS in Figure 5.
Figure 5: ROC Curve of PDS on (HotpotQA from GPT3).
@@ C.4 HaloCheck and SelfCheckNLI
HaloCheck (Elaraby et al., 2023) and SelfCheckNLI (Manakul et al., 2023) are two hallucination detection methods for open-ended generation, which can serve as alternative implementations of PSS. HaloCheck uses the same formula of PPSS as PDS-avg, albeit calculating the average over only a subset of all response pairs: $PSS_{HaloCheck}=mean_{i<j}PPSS(R_{i},R_{j})$. It is important to note that $PPSS(A,B)$ is not symmetric with respect to A and B, introducing a nuanced difference in results. In contrast, SelfCheckNLI adopts a distinct PPSS implementation, leveraging DeBERTa-v3-large (He et al., 2021) finetuned to MNLI (Williams et al., 2018). Additionally, SelfCheckNLI also alters the PSS formulation $PSS_{SelfCheckNLI}=mean_{j}PPSS(R,R_{j})$, where $R$ represents a randomly selected response from the set of highly agreeable answers. This method diverges from the full utilization of all responses by focusing the evaluation on the factuality of a single generated response, as highlighted by Elaraby et al. (2023). This deviation from utilizing the collective information present in all responses leads to a notable performance decrease when replacing our PDS implementation with that of PDS-SelfCheckNLI, as evidenced by the results in Table 3.
@@ Appendix D Combining with verify-and-edit
To edit the rationale $r=s_{1},...,s_{n}$, we first generate one question $q_{i}$ for each sentence $s_{i}$ via prompting. The question $q_{i}$ is to verify the sentence $s_{i}$ in the rationale $r$ and is used as a query to retrieve the relevant context $c_{i}$. $c_{i}$ is composed of the top 3 sentences most similar to query $q_{i}$ ranked by the pre-trained Sentence BERT model (Reimers and Gurevych, 2019) in the retrieved documents. Based on the retrieved context $c_{i}$, the answer for verifying question $q_{i}$ is generated, denoted by $s^{\prime}_{i}$. Finally, we have a new rationale $r^{\prime}=s^{\prime}_{1},...,s^{\prime}_{n}$ that is edited leveraging external knowledge. The prompts for verifying question generation and answering again are provided in Table 17,  18, and  19. Specifically, we use the orical setting from Zhao et al. (2023a) to counteract the influence of retriever performance.
@@ Appendix F Prompts
In our study, we employ modified versions of the original CoT prompts (Wei et al., 2022) for both the GSM8K and StrategyQA datasets. As depicted in Table 22 and 23, the experiments were conducted using setups of 6-shot and 4-shot respectively.
For the Play and Physics, the prompts were meticulously crafted by ourselves. The details of these custom-designed prompts are presented in Table 24 and 25.